princeevee:

What is the mean of the data set? 125, 141, 213, 155, 281

princeevee:

@Vocaloid

Vocaloid:

there are 5 numbers so simply add them up and divide by 5

princeevee:

ok

princeevee:

183?

Vocaloid:

good, that's the answer

princeevee:

1 attachment
princeevee:

i solved that one a while ago, im getting the new question

princeevee:

1 attachment
Vocaloid:

using the table we see that studying with music for more than an hour --> 9 people total number that studied with music --> 9 + 6 so 9/(9+6) --> simplify to get your ans

princeevee:

0.60?

Vocaloid:

good

princeevee:

1 attachment
Vocaloid:

select the graph where the arrow starts at 850 exactly and the arrow stops at year 10

Vocaloid:

actually this is kind of a bad question, the only graph that starts at 850 is C (arrow stops at 9 for some reason?) so I would pick that one

princeevee:

1 attachment
Vocaloid:

notice how f(x) keeps getting bigger and bigger as x increases, and gets smaller and smaller as x decreases which answer choice matches?

princeevee:

b?

Vocaloid:

as x decreases, f(x) is getting smaller and smaller

princeevee:

D then?

Vocaloid:

as x increases, f(x) is getting bigger (increasing)

princeevee:

oh, its A,

Vocaloid:

good, well done

princeevee:

1 attachment
Vocaloid:

when you plug in x = 0 you get 12 which means the graph crosses (0,12) (making A and D the only choices) try plugging in x = 2 into the function to see whether to pick A or D

princeevee:

wait, so which one?

princeevee:

A or D?

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid when you plug in x = 0 you get 12 which means the graph crosses (0,12) (making A and D the only choices) *** try plugging in x = 2 into the function to see whether to pick A or D *** \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

evaluate the function at x = 2.

Vocaloid:

you are given f(x) = 2x^2 - 10x + 12 plug in x = 2 and find f(x) then use the result to determine which graph contains that point

princeevee:

i did that and got D

princeevee:

idk if i did it wrong or anything

Vocaloid:

good, D is correct because x = 2, f(x) = 0 and (2,0) is on the graph D

princeevee:

1 attachment
Vocaloid:

pick two points and calculate (change in y)/(change in x) for example: (0,-1) (3,0) how much does y increase by? how much does x increase by?

princeevee:

i did that and got B

Vocaloid:

what did you get for the change in y?

princeevee:

oh, i havent done that, lemme do that

princeevee:

yeah, i still dont get it

Vocaloid:

y increases from -1 to 0 how much did it increase?

princeevee:

1?

Vocaloid:

good now, x changes from 0 to 3, how much did it increase?

princeevee:

3

Vocaloid:

good, so change in y/change in x = (1/3) making the slope 1/3 which eliminates every answer choice except...?

princeevee:

D?

Vocaloid:

excellent, D is the answer

princeevee:

1 attachment
princeevee:

is would be D too right?

Vocaloid:

over the interval 1 <= x <= 4, x increases from 1 to 4 and y increases from 4 to 13 so change in y/change in x = (13-4)/(4-1) = ?

princeevee:

9/3

Vocaloid:

good, and 9/3 = 3 = your ans

princeevee:

18 of 24 Items The equation y=58−x represents the temperature, y, in degrees Fahrenheit and the time, x, in hours from 6 p.m. to 9 a.m. on a certain day. Let x = 0 represent the temperature at 6 p.m. What is the average rate of change in temperature from 6 p.m. to 9 a.m.? −1°F per hour −2°F per hour 1°F per hour 2°F per hour

Vocaloid:

simply find the slope of y = 58 - x you can re-write this as y = -1x + 58 and compare with y = mx + b to identify what m is

princeevee:

i got B

Vocaloid:

for an equation written in the form y = mx + b, m is the slope. m is the number being multiplied to x. apply this logic and find the slope of y = -1x + 58

princeevee:

i'll try

princeevee:

closest i got was near D, i dont think it's right..

Vocaloid:

try not to look at the answer choices and think about the math if the equation looks like y = mx + b, the number m, which is being multiplied to x, is the slope now, we look at y = -1x + 58 which number is being multiplied to x?

princeevee:

-1

Vocaloid:

good, so that's the slope -1 deg F is your answer

princeevee:

okay

princeevee:

1 attachment
Vocaloid:

just start plugging in x values to see which table matches up for example, when x = -2, what is the value of f(x)?

princeevee:

6?

Vocaloid:

good now find the value of g(x) when x = -2

princeevee:

is it 0

Vocaloid:

good now select the answer choice where f(x) = 6 and g(x) = 0 when x = -2

princeevee:

that would be B?

Vocaloid:

excellent, good work

Vocaloid:

can you open a new question, this one is starting to lag

princeevee:

i can tell